A tradesman marks his goods 25% above the cost price and allows his customers 12% reduction on their bills. What percent profit does he make?


1
Expert's answer
2022-03-11T12:32:50-0500

"P=A(1+0.25)(1-0.12)=1.1A."

Percent profit: "\frac{1.1-1}{1}=0.1=10\%."